# Pre-training Generative Recommender with Multi-Identifier Item Tokenization

### v2

@@ Abstract.
To overcome these limitations, we propose MTGRec, which leverages Multi-identifier item Tokenization to augment token sequence data for Generative Recommender pre-training. Our approach is built upon two core innovations: multi-identifier item tokenization and curriculum recommender pre-training. For multi-identifier item tokenization, we leverage the Residual-Quantized Variational AutoEncoder (RQ-VAE) as the tokenizer backbone and treat model checkpoints from adjacent training epochs as semantically relevant tokenizers. This allows each item to be associated with multiple identifiers, enabling a single user interaction sequence to be converted into several token sequences as different data groups. For curriculum recommender pre-training, we introduce a curriculum learning scheme guided by data influence estimation. Specifically, we estimate the influence of each tokenizer’s data using first-order gradient approximation and dynamically adjust the sampling probability of each data group during recommender pre-training. After pre-training, we fine-tune the model using a single tokenizer to ensure accurate item identification for recommendation. Extensive experiments on three public benchmark datasets demonstrate that MTGRec significantly outperforms both traditional and generative recommendation baselines in terms of effectiveness and scalability.
@@ 3.4. Further Analysis
Figure 4. Performance Comparison w.r.t. Temperature Coefficient on the Instrument and Scientific datasets..
@@ 4.2. Generative Recommendation
Generative recommendation (Rajput et al., 2023; Yin et al., 2024; Chen et al., 2024; Li et al., 2024; Paischer et al., 2024), as an emerging and promising paradigm, has demonstrated superior performance on sequential recommendation tasks compared to traditional recommender systems. In the generative paradigm, each item is indexed with an identifier represented by a list of tokens. This process, known as item tokenization, plays a critical role in generative recommendation. Existing item tokenization methods can be broadly categorized into three groups: heuristic approach, text-based approach, and codebook-based approach. Heuristic approach mainly relies on manually defined rules or techniques, such as time order (Hua et al., 2023), item clustering (Si et al., 2023; Wang et al., 2024c), and matrix decomposition (Petrov and Macdonald, 2023; Hua et al., 2023), to construct item identifiers. While these methods are easy to implement, they often fail to capture the implicit relationships between items, limiting their effectiveness. Text-based approach directly utilizes item attributes, such as title, features, and description as identifiers (Li et al., 2023b; Di Palma, 2023; Harte et al., 2023). These methods are usually designed to leverage the internal knowledge of pre-trained language models to improve recommendation performance. However, they suffer from problems such as inconsistent length, semantic ambiguity, and a lack of collaborative information. In contrast, codebook-based approach (Rajput et al., 2023; Ding et al., 2024; Wang et al., 2024b; Qu et al., 2024) adopt learnable codebooks to quantize item embeddings, thereby constructing fixed-length, semantically rich item identifiers. In addition, several recent studies have focused on enhancing codebook learning to better adapt it to recommendation systems. Notable examples include the introduction of collaborative and diversity regularization (Wang et al., 2024a), as well as alignment between the item tokenizer and the generative recommender (Liu et al., 2024b).

### v3

@@ Abstract.
To overcome these limitations, we propose MTGRec, which leverages Multi-identifier item Tokenization to augment token sequence data for Generative Recommender pre-training. Our approach is built upon two core innovations: multi-identifier item tokenization and curriculum recommender pre-training. For multi-identifier item tokenization, we leverage the Residual-Quantized Variational AutoEncoder (RQ-VAE) as the tokenizer backbone and treat model checkpoints from adjacent training epochs as semantically relevant tokenizers. This allows each item to be associated with multiple identifiers, enabling a single user interaction sequence to be converted into several token sequences as different data groups. For curriculum recommender pre-training, we introduce a curriculum learning scheme guided by data influence estimation. Specifically, we estimate the influence of each tokenizer’s data using first-order gradient approximation and dynamically adjust the sampling probability of each data group during recommender pre-training. After pre-training, we fine-tune the model using a single tokenizer to ensure accurate item identification for recommendation. Extensive experiments on three public benchmark datasets demonstrate that MTGRec significantly outperforms both traditional and generative recommendation baselines in terms of effectiveness and scalability. Our code is available at https://github.com/RUCAIBox/MTGRec.
@@ 3.4. Further Analysis
Figure 4. Performance Comparison w.r.t. Temperature Coefficient on the Instrument and Scientific datasets.
@@ 3.4.5. Applying MTGRec on Other Generative Recommendation Methods
Table 5. Item identifier difference w.r.t. different intervals.
Intervals Instrument Scientific Game
First Any First Any First Any
1 0.39% 13.58% 0.27% 11.4% 0.36% 9.36%
5 0.44% 21.26% 0.58% 22.54% 0.58% 21.22%
10 0.51% 29.75% 0.51% 30.68% 0.57% 30.43%
20 0.75% 44.09% 0.71% 47.33% 0.79% 47.42%
30 0.87% 54.94% 0.85% 58.29% 1.14% 59.95%
3.4.6. Multiple Identifier Difference Analysis
In this section, we analyze the relevance and differences of item identifiers generated by item tokenizers at different training epochs. Specifically, given two sets of item identifiers generated by two item tokenizers, we calculate two metrics: (1) the proportion of items whose first token changed, and (2) the proportion of items with any token changes in the item identifier. The results are shown in Table 5. We observe that for two adjacent tokenizers (i.e., with an epoch interval of 1), the item identifiers on the three datasets exhibit minimal changes, indicating strong semantic consistency. As the epoch interval between tokenizers increases, a larger number of item identifiers change (e.g., 59.95% in the Game dataset), leading to a higher risk of semantic conflict. It is worth noting that even when the interval is large and many identifiers differ, the proportion of changes in the first token typically remains below 1%, thereby preserving the core semantic information.
Table 6. Efficiency of different Methods.
Intervals Instrument Scientific Game
Time Epoch Time Epoch Time Epoch
TIGER 1.33 h 186 1.04 h 184 2.19 h 253
TIGER++ 1.22 h 178 1.02 h 187 2.23 h 264
MTGRec 1.41 h 209 1.21 h 217 2.11 h 248
3.4.7. Efficiency Analysis
In this section, we further investigate the efficiency of the proposed method. As shown in Table 6, we measure the training time and the number of epochs required for model convergence across different methods under the same settings and the same hardware environment. Our method, MTGRec, consists of a pre-training phase of 200 epochs, followed by a low-cost fine-tuning phase. The results demonstrate that our multi-identifier pre-training strategy does not introduce excessive training time costs compared to baseline methods, while achieving significant performance gains. Furthermore, the proposed curriculum learning scheme accelerates model convergence on the Game dataset.
@@ 4.2. Generative Recommendation
Generative recommendation (Rajput et al., 2023; Yin et al., 2024; Chen et al., 2024; Li et al., 2024; Paischer et al., 2024), as an emerging and promising paradigm, has demonstrated superior performance on sequential recommendation tasks compared to traditional recommender systems. In the generative paradigm, each item is indexed with an identifier represented by a list of tokens. This process, known as item tokenization, plays a critical role in generative recommendation. Existing item tokenization methods can be broadly categorized into three groups: heuristic approach, text-based approach, and codebook-based approach. Heuristic approach mainly relies on manually defined rules or techniques, such as time order (Hua et al., 2023), item clustering (Si et al., 2023; Wang et al., 2024c), and matrix decomposition (Petrov and Macdonald, 2023; Hua et al., 2023), to construct item identifiers. While these methods are easy to implement, they often fail to capture the implicit relationships between items, limiting their effectiveness. Text-based approach directly utilizes item attributes, such as title, features, and description as identifiers (Li et al., 2023b; Di Palma, 2023; Harte et al., 2023; Hou et al., 2025). These methods are usually designed to leverage the internal knowledge of pre-trained language models to improve recommendation performance. However, they suffer from problems such as inconsistent length, semantic ambiguity, and a lack of collaborative information. In contrast, codebook-based approach (Rajput et al., 2023; Ding et al., 2024; Wang et al., 2024b; Qu et al., 2024) adopt learnable codebooks to quantize item embeddings, thereby constructing fixed-length, semantically rich item identifiers. In addition, several recent studies have focused on enhancing codebook learning to better adapt it to recommendation systems. Notable examples include the introduction of collaborative and diversity regularization (Wang et al., 2024a), as well as alignment between the item tokenizer and the generative recommender (Liu et al., 2024b).